OpenStudy (austinl):

Solve the given difference equation in terms of the initial condition \(y_0\). Describe the behavior of the solution as \(n\rightarrow\infty\) \(\large{y_{n+1}=\sqrt{\dfrac{n+3}{n+1}}y_n}\)

OpenStudy (austinl):

Um.... what?? I am sorry, missed the lecture for this so I am completely lost.

OpenStudy (atlas):

\[y _{n+1}^2 = \frac{ (n+3)}{( n+1) } y_n^2\]

OpenStudy (atlas):

y1^2/y0^2 = 3 .....(i) y2^2/y1^2 = 2 ......(ii) .......... yn+1^2/yn^2 = n+3/n+1 .... Multiply all of them

OpenStudy (atlas):

\[\frac{ y ^2_{n+1} }{ y ^2_{0} } = \Pi(n+3)/(n+1) \]

OpenStudy (atlas):

the limits of pi are from n=1 to n+1

OpenStudy (atlas):

did you get the answer ??

OpenStudy (austinl):

I'm trying to figure out what you just did...

OpenStudy (atlas):

i m off to sleep .......the answer i guess is (n+2)(n+3) /2.............

OpenStudy (austinl):

Ok, goodnight and thanks for the help!

OpenStudy (atlas):

you are welcome......btw I never took a difference equation class but the trick everytime is to cancel out all other yn's by subtraction or division